# Stealthy Patch-Wise Backdoor Attack in 3D Point Cloud via Curvature Awareness

### v1

@@ Abstract
Backdoor attacks pose a severe threat to deep neural networks (DNN) by implanting hidden backdoors that can be activated with predefined triggers to manipulate model behaviors maliciously. Existing 3D point cloud backdoor attacks primarily rely on sample-wise global modifications, resulting in suboptimal stealthiness. To address this limitation, we propose Stealthy Patch-Wise Backdoor Attack (SPBA), which employs the first patch-wise trigger for 3D point clouds and restricts perturbations to local regions, significantly enhancing stealthiness. Specifically, SPBA decomposes point clouds into local patches and evaluates their geometric complexity using a curvature-based patch imperceptibility score, ensuring that the trigger remains less perceptible to the human eye by strategically applying it across multiple geometrically complex patches with lower visual sensitivity. By leveraging the Graph Fourier Transform (GFT), SPBA optimizes a patch-wise spectral trigger that perturbs the spectral features of selected patches, enhancing attack effectiveness while preserving the global geometric structure of the point cloud. Extensive experiments on ModelNet40 and ShapeNetPart demonstrate that SPBA consistently achieves an attack success rate (ASR) exceeding 96.5% across different models while achieving state-of-the-art imperceptibility compared to existing backdoor attack methods.
@@ 1 Introduction
Point clouds serve as a highly effective data structure for representing 3D geometric data [10], finding increasing utility in safety-critical applications including autonomous driving [14], healthcare [21, 41], and robotics [47]. With this expanding deployment, the security of point cloud models becomes paramount, particularly concerning their vulnerability to backdoor attacks [13]. In these attacks, adversaries inject a pre-defined trigger into a small portion of training data. This manipulation causes the backdoored model, upon training, to consistently misclassify samples containing the trigger into a target class, while maintaining accurate predictions for benign inputs. A well-designed trigger remains inconspicuous within the data yet potently activates malicious behavior in the compromised model. Therefore, trigger design is crucial for both the effectiveness and stealth of backdoor attacks [18, 35, 30].
@@
Figure 1: Comparison of sample-wise and patch-wise trigger injection in 3D point cloud backdoor attacks. Green points represent original samples, while red points indicate modifications introduced by injected triggers.
@@
Early explorations into 3D point cloud backdoor attacks utilized direct global geometric modifications as triggers, such as adding spherical point clusters or applying specific rotations [13, 37, 42]. These global modifications produce deformations readily discernible to the human eye and susceptible to standard data preprocessing techniques like Statistical Outlier Removal (SOR) [45] or rotation augmentation [7]. Recent efforts have shifted towards incorporating local context into trigger design to enhance stealthiness. For instance, IRBA [7] utilized a nonlinear and local transformation to generate sample-wise triggers, while IBAPC [5] employed spectral triggers to balance global shape preservation with localized point perturbations. These triggers are more adaptable, fitting the local geometric structure of samples [7] while preserving the overall global geometry for enhanced stealthiness [5]. However, these methods do not sufficiently focus on local context, limiting their potential for improved concealment. Their triggers either alter the spatial relationship between local and global structures or induce broader deformations, even when aiming for local perturbations.
@@
To this end, we aim to design a novel and highly covert backdoor attack trigger that solely leverages local context information for subtle local modifications. In the 2D image domain, local triggers are typically designed as unified patch-wise triggers [9, 27], which blend seamlessly into diverse image backgrounds. However, the absence of an inherent grid structure in 3D point clouds makes it challenging to design and inject consistent patch-wise deformation triggers without introducing noticeable anomalies or compromising geometric integrity. To address this challenge, we design the patch-wise trigger in the spectral domain rather than the raw spatial domain. By transforming local patches into the spectral domain, we can preserve low-frequency components to maintain the overall local geometric structure while modifying high-frequency components to alter point-level fine-grained details. These modifications are confined within the patch, ensuring precision and subtlety. Furthermore, the patch-wise trigger allows us to selectively perturb specific regions of the point cloud. As noted by [22], human perception varies across different regions of a 3D point cloud, with perturbations in geometrically complex areas being less perceptible. Building on this insight, we selectively filter out smooth patches and apply the patch-wise trigger exclusively to geometrically complex regions, thereby enhancing stealthinessas, as illustrated in Figure 1.
@@
In this paper, we propose the Stealthy Patch-Wise Backdoor Attack (SPBA), which introduces the first patch-wise trigger for 3D point cloud backdoor attack. Our approach first decomposes a point cloud sample into local patches, enabling finer control over trigger injection. To select geometrically complex patches, we design a curvature-based Patch Imperceptibility Score (PIS) to quantify patch complexity. The patches with high PIS are selected and then converted into k-nearest neighbors graphs, providing a structured representation of local geometric relationships. We employ the Graph Fourier Transform (GFT) to map point coordinates into the spectral domain, where perturbations can be introduced in a controlled and structured manner. To maintain the geometric integrity of the affected patches, we apply an optimized spectral trigger that selectively modifies the spectral components while preserving the geometric structure of the local patch. Finally, the Inverse Graph Fourier Transform (IGFT) is utilized to reconstruct the poisoned patches, which are then aggregated with the unmodified regions to form the final poisoned point cloud sample. An alternating optimization process is employed to jointly refine both the point cloud model and the patch-wise trigger, ensuring optimal effectiveness and stealthiness of the trigger. Our main contributions include:
@@
  • We propose SPBA, which introduces the first patch-wise trigger for 3D point cloud backdoor attack based on the Graph Fourier Transform (GFT), enabling precise local spectral perturbations while preserving global geometric structure.
@@
  • We introduce a curvature-based patch imperceptibility scoring mechanism to identify and select geometrically complex patches for trigger injection, significantly enhancing stealthiness by avoiding perturbations in smooth areas.
  • Extensive experiments on ModelNet40 and ShapeNetPart demonstrate the superior attack performance and stealthiness of the proposed SPBA.
@@ 2 Related Work
2.1 Backdoor Attacks in 2D Image
Backdoor attacks pose a significant security threat to deep neural networks (DNNs) by embedding hidden triggers during the training phase, enabling adversaries to manipulate model predictions during inference [17]. The effectiveness of a backdoor attack heavily depends on trigger design, which can be broadly categorized into global triggers and local patch-wise triggers. Prior works have explored global triggers through image generation [15], image warping [23], and frequency-domain injection [6], significantly increasing the difficulty of detection. Additionally, optimization-based approaches [4, 39] refine global triggers using carefully designed objective functions, aiming to balance stealthiness and attack effectiveness. As for local patch-wise triggers, BadNets [9] first demonstrated the feasibility of backdoor attacks in the 2D image domain with a fix patch at the corner. Since then, patch-wise triggers have become a fundamental paradigm in backdoor attacks due to their high attack success rates and stealthiness. Chen  et al. [3] introduced a blending strategy, embedding triggers with partial transparency to make them less detectable. Yuan  et al. [42] found that Vision Transformers (ViTs) are particularly vulnerable to patch-wise triggers, due to their global dependencies within patches. Some works further explored the adaptability of patch-wise trigger to various training paradigms, including self-supervised learning [27] and multimodal contrastive learning [1]. Gao  et al. [8] further improved patch-wise attacks by designing a sparse and invisible trigger through bi-level optimization, minimizing visual detectability while preserving attack effectiveness. What’s more, Li  et al. [16] adopted patch-wise triggers as a harmless and stealthy way to protect the dataset copyright.
However, existing local patch-wise trigger designs primarily focus on 2D images, and their direct adaptation to 3D point clouds presents unique challenges due to the unordered and irregular structure of point cloud data. This motivates our work, which aims to explore optimized local patch-wise triggers for 3D point cloud models.
2.2 Backdoor Attacks in 3D Point Cloud
@@
Recent studies have explored backdoor attacks in 3D point cloud, proposing various trigger designs based on their unique properties [32, 13]. PointBA [13] introduced two attack strategues: PointBA-I, which injects an additional cluster of points as a trigger, and PointBA-O, which exploits a specific rotation angle as the backdoor trigger. Similar to PointBA-I, PCBA [37] embeds a spherical point cloud at an optimized location to manipulate model predictions. To further enhance stealthiness, IRBA [7] achieved sample-specific triggers by utilizing a nonlinear and local transformation. Additionally, iBA [2] generated poisoned samples using a folding-based autoencoder (AE), introducing a sample-wise trigger consists of geometric perturbations during reconstruction. Among spectral-based approaches [11, 12, 20], IBAPC [5] optimized a sample-wise spectral trigger to balance global shape preservation with localized point perturbations. However, the above attack methods all rely on sample-wise trigger injection to generate poisoned samples, which inevitably introduces noticeable global geometric deformations or noise.
@@
In contrast, our approach is the first to propose a local patch-wise trigger for 3D point cloud backdoor attacks, ensuring that perturbations remain confined to selected geometrically complex regions rather than affecting the entire structure. We observe that similar techniques have been employed in recent adversarial attacks [19, 38]. However, our approach differs from these methods in two key aspects: (1) Region selection strategy: Prior adversarial attacks leverage spectral analysis to filter and select different regions for perturbation. In contrast, our method introduces a curvature-based patch imperceptibility score and selects points based on individual patch level. (2) Attack paradigm: As test-time attacks, these adversarial attacks require sample-specific perturbations, which are optimized uniquely for each input in the spatial domain using region-specific optimization. In contrast, our backdoor attack is a train-time attack, requiring a unified patch-wise trigger that can be consistently embedded into local patches across different samples. This need for a local and transferable trigger motivates our spectral-domain design, ensuring that perturbations remain structurally coherent while maintaining high attack effectiveness.
@@ 3.1 Problem Statement
We define a training dataset with $N$ samples as $D={\{(P_{i},Y_{i})\}}^{N}_{i=1}$, where $P_{i}\in\mathbb{R}^{n\times 3}$ represents the three-dimensional coordinates of a point cloud instance and its corresponding label $Y_{i}$ belongs to a predefined set of $K$ classes, i.e., $Y_{i}\in\{1,\dots,K\}$. Building upon the previous explanation [9, 6], we divide the training dataset into two subsets: a benign subset $D_{b}$ and a poisoned subset $D_{p}$. The poisoned subset consists of a portion $\rho$ of the training dataset $D$, where selected samples are manipulated by the trigger injection function $\mathcal{T}$ (as shown in Figure 2) and their labels are modified to a predefined target label $Y_{t}$. Given that the attacker has full control over the training process and access to the entire dataset, the backdoored model $f_{\theta}$ trained on this training dataset maintains high accuracy on benign samples while consistently misclassifying any poisoned sample ${P^{\prime}}_{i}=\mathcal{T}(P_{i})$ as the target class $Y_{t}$.
Figure 2: The framework of our proposed SPBA method. Given a benign point cloud sample $P$, SPBA first decomposes the sample into local patches and evaluates their patch imperceptibility scores (PIS). Patches with high PIS are then selected and transformed into the spectral domain using the Graph Fourier Transform (GFT). A spectral trigger is injected into the selected patches, ensuring stealthy perturbations. Finally, the Inverse Graph Fourier Transform (IGFT) reconstructs the local patches, which are then aggregated with the unmodified patches to form the final poisoned sample $P^{\prime}$.
@@ 3.2 Patchification and Curvature-Aware Selection
Due to the unordered and irregular nature of point clouds, it is infeasible to directly segment patches for trigger injection in the same manner as in 2D images. To address this challenge, we draw inspiration from previous works [46, 43] and employ Farthest Point Sampling (FPS) and K-Nearest Neighbors (KNN) algorithms to systematically decompose each point cloud into local patches. This approach ensures a structured yet flexible partitioning that facilitates effective curvature-aware patch selection for spectral trigger embedding. Given an input point cloud sample $P$, FPS is first applied to select $g$ center points $P^{ct}=\{p^{ct}_{1},p^{ct}_{1},...,p^{ct}_{g}\}$, ensuring a well-distributed set of key reference points. For each center point $p^{ct}\in P^{ct}$, a local patch will be constructed by the KNN algorithm containing $k_{p}$ neighbors: $S_{i}=\mathit{K\!N\!N}(P,p^{ct})$ where $S_{i}\in\mathbb{R}^{k_{p}\times 3}$. Each local patch captures fine-grained geometric information, enabling the assessment of its structural complexity for curvature-aware selection. The resulting set of patches is represented as $\{S_{1},S_{2},...,S_{g}\}$.
As human visual sensitivity varies across different regions of a 3D object, perturbations in high-complexity patches are less perceptible due to the absence of a fixed prior shape. We define a patch imperceptibility score (PIS), which draws from the imperceptibility score concept of the saliency and imperceptibility score (SI score) [22]. However, our PIS is specifically designed to focus on patch-wise geometric complexity without incorporating the saliency score. The local curvature around each point $p_{j}$ is defined as:
$C(p_{j})=\frac{1}{k_{c}}\sum_{q\in\mathcal{N}_{p_{j}}}|\langle\frac{q-p_{j}}{\|q-p_{j}\|_{2}},\mathbf{n_{p_{j}}}\rangle|,$ (1)
where $N_{p_{j}}$ represents the $j$-th patch with $k_{c}$ nearest neighbor of point $p_{j}$ and $\mathbf{n_{p_{j}}}$ denotes its normal vector. To evaluate imperceptibility, the imperceptibility score (IS) of a certain point is computed as the standard deviation of curvature between the point $p_{j}$ and its neighboring points $N_{p_{j}}$, capturing local curvature fluctuations:
$IS(p_{j})=\sqrt{\frac{1}{k_{c}}\sum_{q\in\mathcal{N}_{p_{j}}}(C(q)-\frac{1}{k_{c}}\sum_{q\in\mathcal{N}_{p_{j}}}C(q))^{2}}.$ (2)
Using this measure, we construct an imperceptibility map for the entire point cloud. Finally, the patch imperceptibility score (PIS) of local patch $S$ is calculated by averaging score of all points within the patch:
[…]
From the generated patches, we select the top $m$ patches $\{S^{\prime}_{1},S^{\prime}_{2},...,S^{\prime}_{m}\}$ with the highest PIS for subsequent trigger injection, ensuring that perturbations remain both effective and inconspicuous.

### v2

@@ Abstract
Backdoor attacks pose a severe threat to deep neural networks (DNNs) by implanting hidden backdoors that can be activated with predefined triggers to manipulate model behaviors maliciously. Existing 3D point cloud backdoor attacks primarily rely on sample-wise global modifications, which suffer from low imperceptibility. Although optimization can improve stealthiness, optimizing sample-wise triggers significantly increases computational cost. To address these limitations, we propose the Stealthy Patch-Wise Backdoor Attack (SPBA), the first patch-wise backdoor attack framework for 3D point clouds. Specifically, SPBA decomposes point clouds into local patches and employs a curvature-based imperceptibility score to guide trigger injection into visually less sensitive patches. By optimizing a unified patch-wise trigger that perturbs spectral features of selected patches, SPBA significantly enhances optimization efficiency while maintaining high stealthiness. Extensive experiments on ModelNet40 and ShapeNetPart further demonstrate that SPBA surpasses prior state-of-the-art backdoor attacks in both attack effectiveness and resistance to defense methods.
@@ 1 Introduction
Point clouds serve as a highly effective data structure for representing 3D geometric data [6], finding increasing utility in safety-critical applications including autonomous driving [11], healthcare [15, 32], and robotics [38]. With this expanding deployment, the security of point cloud models becomes paramount, particularly concerning their vulnerability to backdoor attacks [10]. In these attacks, adversaries inject a pre-defined trigger into a small portion of training data. This manipulation causes the backdoored model, upon training, to consistently misclassify samples containing the trigger into a target class, while maintaining accurate predictions for benign inputs. A well-designed trigger remains inconspicuous within the data yet potently activates malicious behavior in the backdoored model. Therefore, trigger design is crucial for both the effectiveness and stealth of backdoor attacks [24, 3, 12, 28].
@@
Figure 1: Comparison of latest sample-wise attacks and our stealthy patch-wise attack. Green points represent the original sample, while red points indicate modifications introduced by injected triggers.
@@
Early explorations into 3D point cloud backdoor attacks utilized direct global geometric modifications as triggers, such as adding spherical point clusters or applying specific rotations [10, 29, 33]. These global modifications produce deformations readily discernible to the human eye and susceptible to standard data preprocessing techniques like Statistical Outlier Removal (SOR) [36] or rotation augmentation [4]. Recent efforts have shifted towards incorporating local context into trigger design to enhance stealthiness. For instance, IRBA [4] utilized a nonlinear and local transformation to generate sample-wise triggers, while IBAPC [2] employed global spectral triggers to balance global shape preservation with localized point perturbations. These triggers are more adaptable, fitting the local geometric structure of samples [4] while preserving the overall global geometry for enhanced stealthiness [2]. However, these methods do not sufficiently focus on local context, limiting their potential for improved stealthiness. Their triggers either alter the spatial relationship between local and global structures or induce broader deformations, even when aiming for local perturbations.
@@
To this end, we aim to design a novel and highly imperceptible backdoor attack trigger that solely leverages local context information for subtle local modifications. In the 2D image domain, local triggers are typically designed as unified patch-wise triggers [5, 21], which blend seamlessly into diverse image backgrounds. However, the absence of an inherent grid structure in 3D point clouds makes it challenging to design and inject consistent patch-wise deformation triggers without introducing noticeable anomalies or compromising geometric integrity. To address this challenge, we design the patch-wise trigger in the spectral domain rather than the raw spatial domain. Specifically, the Graph Fourier Transform (GFT) is employed to map point coordinates of local patches into the spectral domain, where the overall local geometric structure is maintained while the point-level fine-grained details are perturbed. As noted by [16], human perception varies across different regions of a 3D point cloud, with perturbations in geometrically complex areas being less perceptible. Building on this insight, we selectively filter out smooth patches and apply the patch-wise trigger exclusively to geometrically complex regions, thereby enhancing stealthiness, as illustrated in Fig. 1. Notably, the patch-wise spectral trigger requires significantly lower computational cost compared to the sample-wise spectral trigger in the previous IBAPC attack [2]. This greatly improves the efficiency and deployability of our attack method. What’s more, the enriched optimization landscape enabled by patch-wise decomposition, together with targeted patch optimization, is empirically found to be beneficial for improving attack effectiveness.
@@
In this paper, we propose the Stealthy Patch-Wise Backdoor Attack (SPBA), which introduces the first patch-wise trigger in 3D point cloud backdoor attack. Our approach first decomposes a point cloud sample into local patches, enabling finer control over trigger injection. To select geometrically complex patches, we design a curvature-based Patch Imperceptibility Score (PIS) to quantify patch complexity. The patches with high PIS are selected and then converted into KNN graphs, providing a structured representation of local geometric relationships. We employ the Graph Fourier Transform to map point coordinates into the spectral domain, where perturbations can be introduced in a controlled and structured manner. To maintain the geometric integrity of the affected patches, we apply an optimized spectral trigger that selectively modifies the spectral features while preserving the geometric structure. Finally, the Inverse Graph Fourier Transform (IGFT) is utilized to reconstruct the poisoned patches, which are then aggregated with the unmodified regions to form the final poisoned point cloud sample. An alternating optimization process is employed to jointly refine both the point cloud model and the patch-wise trigger, ensuring optimal effectiveness and stealthiness of the trigger. Our main contributions include:
@@
  • We propose SPBA, the first patch-wise backdoor attack on 3D point clouds that leverages the Graph Fourier Transform (GFT) to enable efficient and precise spectral perturbations at the local level, while preserving the global geometric structure.
@@
  • We introduce a curvature-based patch imperceptibility scoring mechanism to identify and select geometrically complex patches for trigger injection, enhancing the stealthiness and effectiveness of SPBA.
  • We conduct comprehensive experiments on two public benchmarks to demonstrate the superiority of SPBA over previous state-of-the-art backdoor attacks on 3D point clouds in terms of attack effectiveness, stealthiness, and efficiency.
@@ 2 Related Work
Recent studies have explored backdoor attacks in 3D point clouds, proposing various trigger designs based on their unique properties [26, 10]. PointBA [10] introduced two attack strategies: PointBA-I, which injects an additional cluster of points as a trigger, and PointBA-O, which exploits a specific rotation angle as the backdoor trigger. Similar to PointBA-I, PCBA [29] embeds a spherical point cloud at an optimized location to manipulate model predictions. To further enhance stealthiness, IRBA [4] achieved sample-specific triggers by utilizing a nonlinear and local transformation. Additionally, iBA [1] generated poisoned samples using a folding-based autoencoder (AE), introducing a sample-wise trigger that consists of geometric perturbations during reconstruction. However, the above attack methods all rely on sample-wise trigger injection to generate poisoned samples, which inevitably introduces noticeable global geometric deformations or noise. To enhance trigger stealthiness, the prior spectral-domain attack method [2] optimized a sample-wise spectral trigger, which caused substantial computational cost and consequently limited both algorithmic efficiency and real-world applicability.
@@
In contrast, our approach is the first to optimize a local patch-wise trigger for 3D point cloud backdoor attacks, achieving favorable results in terms of attack effectiveness, stealthiness, and computational efficiency. We observe that similar techniques have been employed in recent adversarial attacks [13, 30, 7, 8, 14]. However, our approach differs from these methods in two key aspects: (1) Attack paradigm: As test-time attacks, these adversarial attacks require sample-specific perturbations, which are optimized uniquely for each input in the spatial domain using region-specific optimization. In contrast, our backdoor attack is a train-time attack, requiring a unified patch-wise trigger that can be consistently embedded into local patches across different samples. This need for a local and transferable trigger motivates our spectral-domain design, ensuring that perturbations remain structurally coherent while maintaining high attack effectiveness; (2) Region selection strategy: Prior adversarial attacks leverage sample-level spectral analysis to filter and select different regions for perturbation. In contrast, our method introduces a curvature-based patch imperceptibility score and selects points based on the individual patch level.
@@ 3.1 Problem Statement
We define the training dataset with $N$ samples as $D=\{(X_{i},Y_{i})\}_{i=1}^{N}$, where $P_{i}\in\mathbb{R}^{n\times 3}$ denotes the 3D coordinates of the $i^{th}$ sample, and $Y_{i}\in\{1,\dots,K\}$ is its corresponding label from a predefined set of $K$ classes. Building upon the previous explanation [5, 3], we divide the training dataset into two subsets: a benign subset $D_{b}$ and a poisoned subset $D_{p}$. The poisoned subset consists of a portion $\rho$ of the training dataset $D$, where selected samples are manipulated by the trigger injection function $\mathcal{T}$ (as shown in Fig. 2) and their labels are modified to a predefined target label $Y_{t}$. Given that the attacker has full control over the training process and access to the entire dataset, the backdoored model $f_{\theta}$ trained on this training dataset maintains high accuracy on benign samples while misclassifying any poisoned sample $X^{p}_{i}=\mathcal{T}(X_{i})$ as the target class $Y_{t}$.
Figure 2: The framework of our proposed SPBA method. Given a benign point cloud sample $X$, SPBA first decomposes the sample into local patches and evaluates their patch imperceptibility scores (PIS). Patches with high PIS are then selected and transformed into the spectral domain using the Graph Fourier Transform (GFT). A spectral trigger is injected into the selected patches, ensuring stealthy perturbations. Finally, the Inverse Graph Fourier Transform (IGFT) reconstructs the local patches, which are then aggregated with the unmodified patches to form the final poisoned sample $X^{p}$.
@@ 3.2 Patchification and Curvature-Aware Selection
Due to the unordered and irregular nature of point clouds, it is infeasible to directly segment patches for trigger injection in the same manner as in 2D images. To address this challenge, we draw inspiration from previous works [37, 34] and employ Farthest Point Sampling (FPS) and K-Nearest Neighbors (KNN) algorithms to systematically decompose each point cloud into local patches. This strategy enables a structured yet flexible partitioning of the point cloud into local patches. Each patch retains sufficient fine-grained geometric detail, supporting curvature-aware evaluation essential for spectral trigger embedding. Given an input point cloud sample $X$, FPS is first applied to select $g$ center points $X^{ct}=\{x^{ct}_{1},x^{ct}_{1},...,x^{ct}_{g}\}$, ensuring a well-distributed set of key reference points. For each center point $x^{ct}\in X^{ct}$, a local patch will be constructed by the KNN algorithm containing $k_{g}$ neighbors: $S_{i}=\mathit{K\!N\!N}(X,x_{i}^{ct})$ where $S_{i}\in\mathbb{R}^{k_{g}\times 3}$. The resulting set of patches is represented as $\{S_{1},S_{2},...,S_{g}\}$.
As human visual sensitivity varies across different regions of a 3D object, perturbations in high-complexity patches are less perceptible due to the absence of a fixed prior shape. We define a patch imperceptibility score (PIS), which draws from the imperceptibility score concept of the saliency and imperceptibility score (SI score) [16]. However, our PIS is specifically designed to focus on patch-wise geometric complexity without incorporating the saliency score. The local curvature around each point $x_{j}$ in a point cloud $X$ is defined as:
$C(x_{j};X)=\frac{1}{k_{c}}\sum_{q\in\mathcal{N}_{x_{j}}}|\langle\frac{q-x_{j}}{\|q-x_{j}\|_{2}},\mathbf{n_{x_{j}}}\rangle|,$ (1)
where $\mathcal{N}_{x_{j}}$ is defined as the set of $k_{c}$ nearest neighbors of point $x_{j}$ and $\mathbf{n_{x_{j}}}$ denotes its normal vector. The imperceptibility score (IS) of each point is further computed as the standard deviation of curvature between the point $x_{j}$ and its neighboring points $\mathcal{N}_{x_{j}}$, capturing local curvature fluctuations:
$IS(x_{j};X)=\sqrt{\frac{1}{k_{c}}\sum_{q\in\mathcal{N}_{x_{j}}}(C(q)-\frac{1}{k_{c}}\sum_{q\in\mathcal{N}_{x_{j}}}C(q))^{2}}.$ (2)
Using this measure, we construct an imperceptibility map for the entire point cloud. Finally, the patch imperceptibility score (PIS) of decomposed local patch $S_{i}$ is calculated by averaging the score of all points within the patch:
$\mathit{PIS}(S_{i})=\frac{1}{k_{g}}\sum_{x\in{S_{i}}}\mathit{IS}(x).$ (3)
From the decomposed patches, we select the patches $\{S^{\prime}_{1},S^{\prime}_{2},...,S^{\prime}_{m}\}$ with the top-$m$ ranked PIS for subsequent trigger injection, making perturbations imperceptible.